Given that M = 2 −1
−3 4
and that M^2 − 6M + kI = 0, find k.
1
Expert's answer
2020-05-28T18:01:34-0400
"M=\begin{pmatrix}
 2 & -1 \\
 -3 & 4
\end{pmatrix}"

"M^2=\begin{pmatrix}
 2 & -1 \\
 -3 & 4
\end{pmatrix}\begin{pmatrix}
 2 & -1 \\
 -3 & 4
\end{pmatrix}="

"=\begin{pmatrix}
 2(2)-1(-3) & 2(-1)-1(4) \\
 -3(2)+4(-3) & -3(-1)+4(4)
\end{pmatrix}="

"=\begin{pmatrix}
 7 & -6 \\
 -18 & 19
\end{pmatrix}"


"M^2-6M=\begin{pmatrix}
 7 & -6 \\
 -18 & 19
\end{pmatrix}-\begin{pmatrix}
 12 & -6 \\
 -18 & 24
\end{pmatrix}="

"=\begin{pmatrix}
 -5 & 0 \\
 0 & -5
\end{pmatrix}=-5\begin{pmatrix}
 1 & 0 \\
 0 & 1
\end{pmatrix}=-5I"

"k=5"